### v2

# On estimates of trace-norm distance between quantum Gaussian states

Abstract
In the paper [1] estimates for the trace-norm distance between two quantum Gaussian states in terms of the mean vectors and covariance matrices were derived and used to evaluate the number of elements in the $\varepsilon-$net in the set of energy-constrained Gaussian states. In the present paper we obtain different estimates; our proof is based on a fidelity-like quantity which we call states overlap, and is more straightforward leading to estimates which are sometimes even more stringent, especially in the cases of pure or gauge-invariant states. They do not depend on number of modes and hence can be extended to the case of bosonic field with infinite number of modes. These derivations are not aimed to replace the useful inequalities from [1]; they just show an alternative approach to the problem leading to different results. The problem studied in this paper can be considered as a noncommutative analog of estimation of the total variance distance between Gaussian probability distributions in the classical probability theory.
Keywords and phrases: quantum Gaussian state, trace-norm distance, quantum states overlap.
1 Introduction and results
In this paper we obtain estimates for trace-norm distance $\left\|\rho_{1}-\rho_{2}\right\|_{1}$ between two quantum Gaussian states in terms of their parameters – the mean vectors and covariance matrices. To our knowledge, such kind of estimates were first derived in [1], where they were used to evaluate the number of elements in the $\varepsilon-$net in the set of energy-constrained Gaussian states implying that such a state can be efficiently learned via quantum tomography. It was also noticed that such estimates are of an independent interest. Indeed, as an example, they could be useful in investigations concerning characterization of quantum Gaussian observables [2]. The problem studied in the present paper can be also considered as a noncommutative analog of estimation of the total variance distance between Gaussian probability distributions in the classical probability theory [3], [4], [5].
Concerning the method of the proof, the authors of [1] observe: “Initially, one might believe that proving our theorem 3 would be straightforward by bounding the trace distance using the fidelity and leveraging the established formula for the fidelity between Gaussian states [6]. However, this approach turns out to be highly non-trivial due to the complexity of such fidelity formula [6], which makes it challenging to derive a bound based on the norm distance between the first moments and the covariance matrices. Instead, our proof technique directly addresses the trace distance without relying on fidelity. It involves a meticulous analysis based on properties of Gaussian channels and recently demonstrated properties of the energy-constrained diamond norm…”
A hypothetical proof based on the fidelity $\mathrm{Tr}\left|\sqrt{\rho_{1}}\sqrt{\rho_{2}}\right|$ indeed looks to be too difficult. In the present paper, instead of the fidelity we try a simpler quantity $\mathrm{Tr}\sqrt{\rho_{1}}\sqrt{\rho_{2}}$, which we call quantum states overlap, and use the corresponding estimate of the trace-norm distance from our paper [7]. This allows us for a more straightforward proof of the estimates which are sometimes even more stringent, especially in the cases of pure or gauge-invariant states. They do not depend on number of modes and hence can be extended to the case of bosonic field with infinite number of modes. However these derivations are not aimed to replace the useful inequalities from [1] which are adjusted for the purposes of that important paper; they just show a possibility of an alternative approach to the problem leading to different estimates.
[…]
We consider two quantum Gaussian states $\rho_{m_{1},\alpha}$ and $\rho_{m_{2},\beta}$ with the mean vectors $m_{1},m_{2}$ and the covariance matrices $\alpha$, $\beta$ (see sec. 2). We introduce the numbers $a=\|\alpha\|$ and $b=\|\beta\|$. Then our estimates are: for pure Gaussian states
$\left\|\rho_{m_{1},\alpha}-\rho_{{}_{m_{2},\beta}}\right\|_{1}\leq 2\sqrt{8\min\left(a,b\right)\left\|m_{1}-m_{2}\right\|_{2}^{2}+\frac{1}{2}\left\|\alpha-\beta\right\|_{2}^{2}},$ (1)
and for gauge invariant (“passive”) states ($m_{1}=m_{2}=0$, $\alpha$ and $\beta$ commute with the commutation matrix $\Delta$)
$\left\|\rho_{\alpha}-\rho_{{}_{\beta}}\right\|_{1}\leq 4\sqrt{\left(a+b\right)\left\|\alpha-\beta\right\|_{1}-\left\|\alpha-\beta\right\|_{2}^{2}}.$ (2)
For arbitrary Gaussian state we were able to derive the following estimate
$\left\|\rho_{m_{1},\alpha}-\rho_{{}_{m_{2},\beta}}\right\|_{1}\leq$ (3)
$2\sqrt{8\min\left(a,b\right)\left\|m_{1}-m_{2}\right\|_{2}^{2}+\frac{1}{2}\left\|\alpha-\beta\right\|_{2}^{2}+\min\left\{P(a,b),P(b,a)\right\}\left\|\alpha-\beta\right\|_{1}},$
where $P$ is a polynomial with coefficients explicitly given in the proof, see sec. 3. Comparing with the formulas for pure and gauge-covariant states, we conjecture that there might be a more clever way to use the state overlap in the case of general Gaussian states to improve the inequality (3).
The proofs are presented in sec. 3. In the concluding sec. 4 we will compare our estimates with those obtained in [1].
2 The algebra of CCR
We first recall some standard material from the books [8], [9]. Let $(Z,\Delta)$ be a finite-dimensional symplectic vector space with $Z=\mathbb{R}^{2s}$ and the canonical commutation matrix
$\Delta=\mathrm{diag}\left[\begin{array}[]{cc}0&1\\ -1&0\end{array}\right]_{j=1,\dots,s},$ (4)
defining the symplectic form $\Delta(z,z^{\prime})=z^{t}\Delta z^{\prime}.$ In what follows the Hilbert space $\mathcal{H}$ is the space of an irreducible representation $z\rightarrow W(z);\,z\in Z,$ of the Weyl canonical commutation relations
$W(z)W(z^{\prime})=\exp[-\frac{i}{2}\Delta(z,z^{\prime})]\,W(z+z^{\prime}),$ (5)
where $W(z)=\exp i\,Rz$ are the unitary operators with the generators
$Rz=\sum_{j=1}^{s}(x_{j}q_{j}+y_{j}p_{j}),\quad z=[x_{j}\,\,y_{j}]_{j=1,\dots,s}^{t}$ (6)
and $R=\left[q_{1},p_{1},\dots,q_{s},p_{s}\right]$ is the row vector of the bosonic position-momentum observables, satisfying the Heisenberg commutation relations on an appropriate domain.
The quantum Fourier transform of a trace class operator $\rho$ is defined as $f(z)=\mathop{\rm Tr}\nolimits\rho\,W(z)$. The quantum Parceval formula holds [8]:
$\mathop{\rm Tr}\nolimits\rho^{\ast}\sigma=\int\,\overline{\mathop{\rm Tr}\nolimits\rho W(z)}\mathop{\rm Tr}\nolimits\sigma W(z)\frac{d^{2s}z}{\left(2\pi\right)^{s}}.$ (7)
The covariance matrix of a state with finite second moments is defined as
$\alpha=\mathrm{Re}\,\mathrm{Tr}\left(R-m\right)^{t}\rho\left(R-m\right),$
where $m=\mathrm{Tr}\rho R$ is the row-vector of the first moments (the mean vector). Matrix $\alpha$ is a real symmetric $2s\times 2s$-matrix satisfying
$\alpha\geq\pm\frac{i}{2}\Delta.$ (8)
The condition (8) implies that matrix $\alpha$ is positive definite. There is a theorem in linear algebra saying that a positive definite matrix can be represented as
$\alpha=\,S^{t}\mathrm{diag}\left[\begin{array}[]{cc}a_{j}&0\\ 0&a_{j}\end{array}\right]_{j==1,\dots,s}S,$
where $S$ is a symplectic matrix ($S^{t}\Delta S=\Delta$) and $a_{j}>0$ (see e.g. [8]). For a positive definite matrix $\alpha$ the inequality
$\alpha+\frac{1}{4}\Delta\alpha^{-1}\Delta\geq 0.$ (9)
is equivalent to (8), because both (8) and (9) express the fact that symplectic eigenvalues of matrix $\alpha$ satisfy $a_{j}\geq 1/2$, $j=1,\dots,s.$
A Gaussian state $\rho_{m,\alpha}$ is determined by its quantum characteristic function
$\mathrm{Tr}\,\rho_{m,\alpha}W(z)=\exp\left(im^{t}z-\frac{1}{2}z^{t}\alpha z\right).$ (10)
Here $\alpha$ is the covariance matrix and $m$ is the mean vector of the state. For a centered state $(m=0)$ we denote $\rho_{\alpha}=\rho_{0,\alpha}.$
The equality
$\alpha+\frac{1}{4}\Delta\alpha^{-1}\Delta=0$ (11)
in (9) holds if and only if $\rho_{m,\alpha}$ is a pure Gaussian state $(a_{j}=1/2;j=1,\dots,s)$.
Let $\alpha$ be a covariance matrix of a Gaussian state, that is a real symmetric matrix satisfying the inequality (8). It is thus positive definite, hence defines the inner product $\alpha(z,z^{\prime})=z^{t}\alpha z^{\prime}$ in $Z$ making it the real Hilbert space $(Z,\alpha).$ If $T$ is an operator in $(Z,\alpha)$, then its adjoint is $\left(T\right)_{\alpha}^{\ast}=\alpha^{-1}T^{t}\alpha.$ We call $T$ $\alpha-$symmetric ($\alpha-$skew-symmetric) if $T=\left(T\right)_{\alpha}^{\ast}$ ($T=-\left(T\right)_{\alpha}^{\ast}$), which amounts to $\alpha T=T^{t}\alpha$ (resp. $\alpha T=-T^{t}\alpha$). An $\alpha-$symmetric operator $T$ is $\alpha-$positive if $\ \alpha(z,Tz)\geq 0$ for all $z,$ which amounts to $\alpha T\geq 0.$ For such an operator the square root $\sqrt{T}$ is the unique $\alpha-$positive operator satisfying $\left(\sqrt{T}\right)^{2}=T.$
Consider the operator [^1] In the case (4) we have $\Delta^{-1}=-\Delta$, but we retain the notation $\Delta^{-1}$ having in mind more general situation where $\Delta$ is an arbitrary nondegenerate skew-symmetrc matrix. Our derivations then remain valid with appropriate modifications. $A=\Delta^{-1}\alpha$ which represents the bilinear form $\alpha(z,z^{\prime})=z^{t}\alpha z^{\prime}$ in the symplectic space $(Z,\Delta):$
$\alpha(z,z^{\prime})=\Delta(z,Az^{\prime}),\quad z,z^{\prime}\in Z.$
This operator is skew-symmetric in the Euclidean space $(Z,\alpha)$, hence $I_{2s}+\left(2\Delta^{-1}\alpha\right)^{-2}$ is $\alpha-$symmetric. Moreover it is $\alpha-$positive as
$\alpha\left[I_{2s}+\left(2\Delta^{-1}\alpha\right)^{-2}\right]=\alpha+\frac{1}{4}\Delta\alpha^{-1}\Delta\geq 0,$ (12)
due to (9). Hence the square root
$\Upsilon_{\alpha}=\sqrt{I_{2s}+\left(2\Delta^{-1}\alpha\right)^{-2}}$ (13)
is defined and satisfies
$\alpha\Upsilon_{\alpha}=\Upsilon_{\alpha}^{t}\alpha\geq 0.$ (14)
Later we will need the following important matrix
$\hat{\alpha}\equiv\alpha\left(I_{2s}+\Upsilon_{\alpha}\right)=\alpha\left(I_{2s}+\sqrt{I_{2s}+\left(2\Delta^{-1}\alpha\right)^{-2}}\right).$ (15)
The matrix $\hat{\alpha}$ is $\alpha-$symmetric and satisfies
$-\frac{1}{4}\Delta\hat{\alpha}^{-1}\Delta=2\alpha-\hat{\alpha}.$ (16)
Indeed by using (15) and noticing that $\left(I_{2s}+\Upsilon_{\alpha}\right)\left(I_{2s}-\Upsilon_{\alpha}\right)=-\left(2\Delta^{-1}\alpha\right)^{-2}$ we have
$\hat{\alpha}^{-1}$ $=$ $-4\Delta^{-1}\alpha\left(I_{2s}-\Upsilon_{\alpha}\right)\Delta^{-1}$ (17)
$=$ $-4\Delta^{-1}\alpha\Delta^{-1}+4\Delta^{-1}\alpha\Upsilon_{\alpha}\Delta^{-1}$ (18)
$=$ $-4\Delta^{-1}\left(2\alpha-\hat{\alpha}\right)\Delta^{-1},$ (19)
where we also used the fact that $\Upsilon_{\alpha}$ is a function of $\Delta^{-1}\alpha$ and hence commutes with it. But this is equivalent to (16).
In particular, (14) and (16) imply
$\alpha\leq\hat{\alpha}\leq 2\alpha,$ (20)
because
$0\leq\frac{1}{4}\Delta^{t}\hat{\alpha}^{-1}\Delta=-\frac{1}{4}\Delta\hat{\alpha}^{-1}\Delta.$ (21)
Also from (17)
$\alpha\left(I_{2s}-\Upsilon_{\alpha}\right)=-\frac{1}{4}\Delta\hat{\alpha}^{-1}\Delta\geq 0$
hence
$\alpha\Upsilon_{\alpha}\leq\alpha.$ (22)
Introducing the number $a=\left\|\alpha\right\|,$ we therefore have the estimates
$a=\left\|\alpha\right\|\leq\left\|\hat{\alpha}\right\|\leq 2\left\|\alpha\right\|=2a.$ (23)
From (9) $\alpha^{-1}\leq-4\Delta\alpha\Delta=4\Delta^{t}\alpha\Delta$ hence taking into account that $\left\|\Delta\right\|=1,$
$a^{-1}=\left\|\alpha\right\|^{-1}\leq\left\|\alpha^{-1}\right\|\leq 4\left\|\alpha\right\|=4a.$ (24)
By (20), $\frac{1}{2}\alpha^{-1}\leq\hat{\alpha}^{-1}\leq\alpha^{-1},$ hence
$\frac{1}{2a}\leq\frac{1}{2}\left\|\alpha^{-1}\right\|\leq\left\|\hat{\alpha}^{-1}\right\|\leq\left\|\alpha^{-1}\right\|\leq 4a.$ (25)
3
Proofs
Our approach is based on the second inequality in
$2\left(1-\mathrm{Tr}\sqrt{\rho_{1}}\sqrt{\rho_{2}}\right)\leq\left\|\rho_{1}-\rho_{2}\right\|_{1}\leq 2\sqrt{1-\left(\mathrm{Tr}\sqrt{\rho_{1}}\sqrt{\rho_{2}}\right)^{2}},$ (26)
where $\rho_{1},\rho_{2}$ are density operators. If the corresponding states are pure then the second inequality turns into equality. The proof given in [7] is partly based on estimates from [10].
The unitary displacement operators $D(z)=W(-\Delta^{-1}z)$ satisfy the equation that follows from the canonical commutation relations (5)
$D(z)^{\ast}W(w)D(z)=\exp\left(iw^{t}z\right)W(w),$ (27)
implying
$\rho_{m_{1},\alpha}=D(m_{1})\rho_{\alpha}D(m_{1})^{\ast},\quad\rho_{m_{2},\beta}=D(m_{1})\rho_{{}_{m,\beta}}D(m_{1})^{\ast},$
where $m=m_{2}-m_{1}.$ Hence
$\left\|\rho_{m_{1},\alpha}-\rho_{m_{2},\beta}\right\|_{1}=\left\|\rho_{\alpha}-\rho_{{}_{m,\beta}}\right\|_{1},$
and in what folows we will deal with the second quantity.
From [7] we know that the quantum Fourier transform of $\sqrt{\rho_{m,\alpha}}$ is
$\mathrm{Tr}\,\sqrt{\rho_{m,\alpha}}W(z)=\sqrt[4]{\det\left(2\hat{\alpha}\right)}\exp\left(im^{t}z-\frac{1}{2}z^{t}\hat{\alpha}z\right),$ (28)
where $\hat{\alpha}$ is given by (15).
Applying the Parceval relation (7) we arrive at the formula
$\mathrm{Tr}\sqrt{\rho_{\alpha}}\sqrt{\rho_{m,\beta}}=\frac{\left(\det\hat{\alpha}\det\hat{\beta}\right)^{1/4}}{\det\sigma(\hat{\alpha},\hat{\beta})^{1/2}}\exp\left(-\frac{1}{2}m^{t}\sigma(\hat{\alpha},\hat{\beta})^{-1}m\right),$ (29)
$\sigma(\hat{\alpha},\hat{\beta})=\frac{\hat{\alpha}+\hat{\beta}}{2}.$
Let us evaluate the factor before the exponent:
$\frac{\left(\det\hat{\alpha}\det\hat{\beta}\right)^{1/4}}{\det\sigma(\hat{\alpha},\hat{\beta})^{1/2}}$
$=$ $\left[\det\left(\frac{1+\hat{\alpha}^{-1}\hat{\beta}}{2}\right)\det\left(\frac{1+\hat{\beta}^{-1}\hat{\alpha}}{2}\right)\right]^{-1/4}$
$=$ $\left[\det\left(\frac{T+2+T^{-1}}{4}\right)\right]^{-1/4},$
where $T=\hat{\alpha}^{-1}\hat{\beta}$ is positive operator in the Hilbert space $Z_{\hat{\alpha}}.$ Also $T-2+T^{-1}$ is positive operator in $Z_{\hat{\alpha}},$ therefore
$\left[\det\left(\frac{T+2+T^{-1}}{4}\right)\right]^{-1/4}$
$=$ $\exp\left[-\frac{1}{4}\mathrm{Tr}\ln\left(\frac{T+2+T^{-1}}{4}\right)\right]$
$=$ $\exp\left[-\frac{1}{4}\mathrm{Tr}\ln\left(I+\frac{T-2+T^{-1}}{4}\right)\right]$
$\geq$ $\exp\left[-\frac{1}{4}\mathrm{Tr}\left(\frac{T-2+T^{-1}}{4}\right)\right],$
where the last inequality follows from $\ln\left(1+x\right)\leq x$ for $x\geq 0.$
Then from (29) it follows
$1-\left(\mathrm{Tr}\sqrt{\rho_{\alpha}}\sqrt{\rho_{m,\beta}}\right)^{2}$
$\leq$ $m^{t}\sigma(\hat{\alpha},\hat{\beta})^{-1}m+\frac{1}{8}\mathrm{Tr}\left(T-2+T^{-1}\right)$
$=$ $2m^{t}\left(\hat{\alpha}+\hat{\beta}\right)^{-1}m+\frac{1}{8}\mathrm{Tr}\left(\hat{\alpha}-\hat{\beta}\right)\left(\hat{\beta}^{-1}-\hat{\alpha}^{-1}\right).$
The second inequality in (26) then implies
$\left\|\rho_{\alpha}-\rho_{{}_{m,\beta}}\right\|_{1}\leq 2\sqrt{2m^{t}\left(\hat{\alpha}+\hat{\beta}\right)^{-1}m+\frac{1}{8}\mathrm{Tr}\left(\hat{\alpha}-\hat{\beta}\right)\left(\hat{\beta}^{-1}-\hat{\alpha}^{-1}\right)}.$ (30)
Let us evaluate the expression $\mathrm{Tr}\left(\hat{\alpha}-\hat{\beta}\right)\left(\hat{\beta}^{-1}-\hat{\alpha}^{-1}\right).$
From (18)
$\hat{\alpha}^{-1}=-4\Delta^{-1}\alpha\Delta^{-1}+4\Delta^{-1}\alpha\Upsilon_{\alpha}\Delta^{-1}.$
By using this and (15) we have
$\mathrm{Tr}\left(\hat{\alpha}-\hat{\beta}\right)\left(\hat{\beta}^{-1}-\hat{\alpha}^{-1}\right)$
$=$ $4\mathrm{Tr}[\left(\alpha-\beta\right)+\left(\alpha\Upsilon_{\alpha}-\beta\Upsilon_{\beta}\right)]\Delta^{-1}[\left(\alpha-\beta\right)-\left(\alpha\Upsilon_{\alpha}-\beta\Upsilon_{\beta}\right)]\Delta^{-1}$
$=$ $4\mathrm{Tr}[\left(\alpha-\beta\right)\Delta^{-1}]^{2}-4\mathrm{Tr}[\left(\alpha\Upsilon_{\alpha}-\beta\Upsilon_{\beta}\right)\Delta^{-1}]^{2}$ (32)
$\leq$ $4\mathrm{Tr}[\alpha-\beta]^{2}+4\mathrm{Tr}[\alpha\Upsilon_{\alpha}-\beta\Upsilon_{\beta}]^{2},$ (33)
where in (33) we used the Cauchy-Schwarz inequality for trace and the fact that $\left\|\Delta^{-1}\right\|=1.$
3.1 Pure states
In the case of pure state the relation (11) implies $\left(2\Delta^{-1}\alpha\right)^{2}=-I_{2s},$ that is $\Upsilon_{\alpha}=0$ and $\hat{\alpha}=\alpha.$ Similarly $\hat{\beta}=\beta$ and $\Upsilon_{\beta}=0$. Then (33) with (30) imply
$\left\|\rho_{\alpha}-\rho_{{}_{m,\beta}}\right\|_{1}\leq 2\sqrt{2m^{t}\left(\alpha+\beta\right)^{-1}m+\frac{1}{2}\mathrm{Tr}\left(\alpha-\beta\right)^{2}}.$
Since $\left(\alpha+\beta\right)^{-1}\leq\alpha^{-1},\left(\alpha+\beta\right)^{-1}\leq\beta^{-1},$ we have by using (25)
$m^{t}\left(\alpha+\beta\right)^{-1}m$$\leq$ $\min\left(\left\|\alpha^{-1}\right\|,\left\|\beta^{-1}\right\|\right)\left\|m\right\|_{2}^{2}$ (34)
$\leq$ $4\min\left(a,b\right)\left\|m\right\|_{2}^{2},$
where $a=\left\|\alpha\right\|,b=\left\|\beta\right\|.$ Thus for pure Gaussian states we obtain
$\left\|\rho_{\alpha}-\rho_{{}_{m,\beta}}\right\|_{1}\leq 2\sqrt{8\min\left(a,b\right)\left\|m\right\|_{2}^{2}+\frac{1}{2}\left\|\alpha-\beta\right\|_{2}^{2}},$ (35)
that is our first estimate (1).
3.2 Gauge-invariant states
Operator $J$ in $(Z,\Delta)$ is called operator of complex structure if $J^{2}=-I_{2s},$ where $I_{2s}$ is the identity operator in $Z$, and it is $\Delta-$positive in the sense that
$\Delta J=-J^{t}\Delta,\quad\Delta J\geq 0.$ (36)
A distinguished complex structure is given by $J_{\Delta}=\Delta^{-1}.$ For any complex structure $J$ we have $J=S^{t}J_{\Delta}S,$ where $S$ is an appropriate symplectic transformation.
The Gaussian state is gauge-invariant with respect to $J$ if the mean vector $m=0$ and the operator $A=\Delta^{-1}\alpha$ satisfies $[A,J]=0$ which is equivalent to $\alpha J=-J^{t}\alpha.$ In the case $J=J_{\Delta}=\Delta^{-1}$ this amounts to $[\alpha,\Delta]=0.$
Since all the complex structures are symplectically equivalent to $J_{\Delta}=\Delta^{-1},$ we can reduce to the case $J=J_{\Delta}$. Let us apply the expression (32) in this case. Then $\Upsilon_{\alpha}=\sqrt{I_{2s}-\left(2\alpha\right)^{-2}}$ and $\hat{\alpha}=\alpha+\sqrt{\alpha^{2}-1/4}$ and similarly for $\beta.$ Taking into account that $J_{\Delta}$ commutes with $\alpha$ and $\beta$ and $J_{\Delta}^{2}=-I_{2s}$, the expression (32) turns into
$-4\mathrm{Tr}[\alpha-\beta]^{2}+4\mathrm{Tr}[\sqrt{\alpha^{2}-1/4}-\sqrt{\beta^{2}-1/4}]^{2}.$ (37)
For the second term, the operators $T_{1}=\sqrt{\alpha^{2}-1/4},T_{2}=\sqrt{\beta^{2}-1/4}$ are positive in the real Hilbert space $(Z,I)$ equipped with inner product $(z,z^{\prime})=z^{t}z^{\prime}$ (corresponding to $\alpha=I_{2s}$). Hence we can use the inequality (1.2) from [7] (see also [10])
$\left\|T_{1}-T_{2}\right\|_{2}^{2}\leq\left\|T_{1}^{2}-T_{2}^{2}\right\|_{1}$ (38)
implying that the second term in r.h.s. of (37) is upperbounded by
$4\left\|\alpha^{2}-\beta^{2}\right\|_{1}$ $=$ $4\left\|\left(\alpha-\beta\right)\circ\left(\alpha+\beta\right)\right\|_{1}$
$\leq$ $4\left(\left\|\alpha\right\|+\left\|\beta\right\|\right)\left\|\alpha-\beta\right\|_{1}$
$=$ $4\left(a+b\right)\left\|\alpha-\beta\right\|_{1},$
where $\circ$ denoted the Jordan product. Inserting in (37) we obtain the final estimate (2) for the trace-norm distance of gauge-invariant states.
3.3 “Brute force” estimate for arbitrary Gaussian states
Consider the real Hilbert space $(Z,\alpha).$ We wish to compare Schatten norms of operators on $(Z,\alpha)$ (called $\alpha-$norms) with norms of the same operators in the Hilbert space $(Z,I)$ equipped with the inner product $z^{t}z^{\prime}$. Recall that the adjoint in $(Z,\alpha)$ is $\left(T\right)_{\alpha}^{\ast}=\alpha^{-1}T^{t}\alpha,$ hence the modulus of the operator is $\left|T\right|_{\alpha}=\sqrt{\alpha^{-1}T^{t}\alpha T}.$ Therefore Schatten $\alpha-$norms for $p\geq 1$ are
$\left\|T\right\|_{p,\alpha}=\left[\mathrm{Tr}\left(\alpha^{-1}T^{t}\alpha T\right)^{p/2}\right]^{1/p}.$
But
$\left(\alpha^{-1}T^{t}\alpha T\right)^{p/2}$ $=$ $\left[\alpha^{-1/2}\left(\alpha^{-1/2}T^{t}\alpha^{1/2}\alpha^{1/2}T\alpha^{-1/2}\right)\alpha^{1/2}\right]^{p/2}$
$=$ $\alpha^{-1/2}\left[\alpha^{-1/2}T^{t}\alpha^{1/2}\alpha^{1/2}T\alpha^{-1/2}\right]^{p/2}\alpha^{1/2}.$
Taking trace, we obtain
$\left\|T\right\|_{p,\alpha}=\left\|\alpha^{1/2}T\alpha^{-1/2}\right\|_{p},$
where $\left\|\cdot\right\|_{p}$ are the norms of operators in $(Z,I)$. This relation is valid also for $p=\infty.$ By using the Hölder inequality for the Schatten norms, we obtain from (24)
$\left\|T\right\|_{p,\alpha}\leq\sqrt{\left\|\alpha\right\|\left\|\alpha^{-1}\right\|}\left\|T\right\|_{p}\leq 2a\left\|T\right\|_{p}$
and similarly $\left\|T\right\|_{p}\leq 2a\left\|T\right\|_{p,\alpha}.$ Hence
$\frac{1}{2a}\left\|T\right\|_{p}\leq\left\|T\right\|_{p,\alpha}\leq 2a\left\|T\right\|_{p}.$ (39)
In the case of the operator $\alpha-$norm which formally corresponds to $p=\infty$ we have similar estimates.
We now come back to our main problem and to the quantity (3) for arbitrary states: according to (33)
$\frac{1}{8}\mathrm{Tr}\left(\hat{\alpha}-\hat{\beta}\right)\left(\hat{\beta}^{-1}-\hat{\alpha}^{-1}\right)\leq\frac{1}{2}\left\|\alpha-\beta\right\|_{2}^{2}+\frac{1}{2}\left\|\alpha\Upsilon_{\alpha}-\beta\Upsilon_{\beta}\right\|_{2}^{2}.$ (40)
Introducing $A=2\Delta^{-1}\alpha,$ $B=2\Delta^{-1}\beta$ (this time we include factor 2 for notational convenience), the second term is evaluated as
$\frac{1}{2}\left\|\alpha\Upsilon_{\alpha}-\beta\Upsilon_{\beta}\right\|_{2}^{2}$$\leq$ $\frac{a^{2}}{2}\left\|\Upsilon_{\alpha}-\alpha^{-1}\beta\Upsilon_{\beta}\right\|_{2}^{2}$
$\leq$ $2a^{4}\left\|\sqrt{I_{2s}+A^{-2}}-A^{-1}B\sqrt{I_{2s}+B^{-2}}\right\|_{2,\alpha}^{2},$
where in the second inequality we used (39). By (14) both operators $\sqrt{I_{2s}+A^{-2}}$ and $A^{-1}B\sqrt{I_{2s}+B^{-2}}=\alpha^{-1}\beta\sqrt{I_{2s}+B^{-2}}$ are Hermitian positive in $Z_{\alpha}.$ Hence we can again use the inequality (38)
$\left\|T_{1}-T_{2}\right\|_{2,\alpha}^{2}\leq\left\|T_{1}^{2}-T_{2}^{2}\right\|_{1,\alpha}$
valid for $\alpha-$positive $T_{1},T_{2},$ obtaining
$2a^{4}\left\|\sqrt{I_{2s}+A^{-2}}-A^{-1}B\sqrt{I_{2s}+B^{-2}}\right\|_{2,\alpha}^{2}$
$\leq$ $2a^{4}\left\|I_{2s}+A^{-2}-\left(A^{-1}B\sqrt{I_{2s}+B^{-2}}\right)^{2}\right\|_{1,\alpha}$
$\leq$ $4a^{5}\left\|I_{2s}+A^{-2}-\left(A^{-1}B\sqrt{I_{2s}+B^{-2}}\right)^{2}\right\|_{1}.$
Expression under the norm is equal to
$I_{2s}+A^{-2}-\left(\sqrt{I_{2s}+B^{-2}}+\left(A^{-1}B-I_{2s}\right)\sqrt{I_{2s}+B^{-2}}\right)^{2}$
$=$ $\left(A^{-2}-B^{-2}\right)$
-- $2\left[\left(A^{-1}B-I_{2s}\right)\Upsilon_{\beta}\right]\circ\Upsilon_{\beta}$
-- $\left(\left(A^{-1}B-I_{2s}\right)\Upsilon_{\beta}\right)^{2}.$
Let us evaluate trace norm of the first term of the r.h.s. By using (25) several times, we obtain
$\left\|A^{-2}-B^{-2}\right\|_{1}$ $=$ $\left\|\left(A^{-1}-B^{-1}\right)\circ\left(A^{-1}+B^{-1}\right)\right\|_{1}$
$\leq$ $\frac{\left\|\alpha^{-1}\right\|+\left\|\beta^{-1}\right\|}{2}\left\|\left(A^{-1}-B^{-1}\right)\right\|_{1}$
$\leq$ $\frac{\left\|\alpha^{-1}\right\|+\left\|\beta^{-1}\right\|}{2}\left\|\beta^{-1}\right\|\left\|A^{-1}B-I_{2s}\right\|_{1}$
$\leq$ $8\left(a+b\right)b\left\|A^{-1}B-I_{2s}\right\|_{1},$
while by (25)
$\left\|A^{-1}B-I_{2s}\right\|_{1}=\left\|\alpha^{-1}\beta-I_{2s}\right\|_{1}\leq 4a\left\|\beta-\alpha\right\|_{1}.$ (41)
Thus
$\left\|A^{-2}-B^{-2}\right\|_{1}\leq 32(a+b)ab\left\|\beta-\alpha\right\|_{1}.$ (42)
For the second term we use the fact that by (22)
$\left\|\Upsilon_{\beta}\right\|_{\beta}=\sup\frac{z^{t}\beta\Upsilon_{\beta}z}{z^{t}\beta z}\leq 1.$ (43)
Then by (39)
$\left\|\left[2\left(A^{-1}B-I_{2s}\right)\Upsilon_{\beta}\right]\circ\Upsilon_{\beta}\right\|_{1}$
$\leq$ $2b\left\|\left[2\left(A^{-1}B-I_{2s}\right)\Upsilon_{\beta}\right]\circ\Upsilon_{\beta}\right\|_{1,\beta}$
$\leq$ $4b\left\|A^{-1}B-I_{2s}\right\|_{1,\beta}.$
By (41) this is evaluated as
$\leq$ $8b^{2}\left\|\alpha^{-1}\beta-I_{2s}\right\|_{1}$
$\leq$ $32ab^{2}\left\|\beta-\alpha\right\|_{1}.$
For the third term, again using (39) and (43),
$\left\|\left(\left(A^{-1}B-I_{2s}\right)\Upsilon_{\beta}\right)^{2}\right\|_{1}$
$\leq$ $2b\left\|\left(\left(A^{-1}B-I_{2s}\right)\Upsilon_{\beta}\right)^{2}\right\|_{1,\beta}$
$\leq$ $2b\left\|A^{-1}B-I_{2s}\right\|_{1,\beta}\left\|A^{-1}B-I_{2s}\right\|_{\beta}$
$\leq$ $8b^{3}\left\|A^{-1}B-I_{2s}\right\|_{1}\left\|A^{-1}B-I_{2s}\right\|.$
From (25) and (41) this is evaluated as
$\leq$ $8b^{3}4a\left\|\beta-\alpha\right\|_{1}4a\left\|\beta-\alpha\right\|$
$=$ $128a^{2}b^{3}\left\|\beta-\alpha\right\|_{1}\left\|\beta-\alpha\right\|$
$\leq$ $128a^{2}b^{3}\left(a+b\right)\left\|\beta-\alpha\right\|_{1}.$
Summarizing
$\frac{1}{2}\left\|\alpha\Upsilon_{\alpha}-\beta\Upsilon_{\beta}\right\|_{2}^{2}$$\leq$ $4a^{5}\left[32(a+b)ab+32ab^{2}+128a^{2}b^{3}\left(a+b\right)\right]\left\|\beta-\alpha\right\|_{1}$
$=$ $P(a,b)\left\|\beta-\alpha\right\|_{1}$
By using (40) we obtain the estimate (3) for the trace-norm distance of two Gaussian states.
4 Discussion
The estimate given in [1] (theorem 2) reads in our notations
$\left\|\rho_{m_{1},\alpha}-\rho_{{}_{m_{2},\beta}}\right\|_{1}\leq\sqrt{2}\left(\sqrt{N}+\sqrt{N+1}\right)\left[\left\|m_{1}-m_{2}\right\|_{2}+2\sqrt{\left\|\alpha-\beta\right\|_{1}}\right].$
The parameter $N$ here has important physical meaning as the total mean photon number (over all modes). Mathematically, it is expressed via the trace norm of the covariance matrices, $N=\max\left(N_{m_{1},\alpha},N_{{}_{m_{2},\beta}}\right),$ $N_{m,\alpha}=\frac{1}{2}\left(\mathrm{Tr}\left[\alpha-\frac{1}{2}\right]+\left\|m\right\|^{2}\right)=\frac{1}{2}\left(\left\|\alpha\right\|_{1}+\left\|m\right\|^{2}-s\right)].$ It is thoroughly adapted to the problem of learning the energy- constrained Gaussian state. For pure states there is improved version (lemma 4):
$\left\|\rho_{m_{1},\alpha}-\rho_{{}_{m_{2},\beta}}\right\|_{1}\leq 2\left(\sqrt{N+s/2}\right)\sqrt{2\left\|m_{1}-m_{2}\right\|_{2}^{2}+2\left\|\alpha-\beta\right\|}.$
One can easily construct photon number distributions over the modes for which $N_{m,\alpha}\,\sim\left\|\alpha\right\|_{1}\gg\left\|\alpha\right\|=a$ and similarly for $\beta.$ In such cases our estimates may be more stringent, notably for pure and gauge-covariant states. Our estimates are also have the advantage of being dimension independent, hence in principle transferrable to the case of infinite dimensions. In that case $\alpha$ and $\beta$ are bounded operators on one-particle Hilbert space, so the constants $a$ and $b$ entering in the coefficients of our estimates are finite while $N$ can be easily infinite.
In the paper [1] also the lower bounds in terms of first and second moments were derived for the trace-norm distance of two Gaussian states (theorem 5). It would be interesting to investigate whether the first inequality in (26) could be used for a similar purpose.
Acknowledgment. The author is grateful to M.E.Shirokov for comments helping to improve the presentation.